OpenStudy (anonymous):

Find the value of y when x equals 14 in this equation. -6x + 6y = -42

OpenStudy (sriram):

y=7

OpenStudy (diyadiya):

-84+6y=-42 6y=42 y=7

OpenStudy (anonymous):

y=42/6=7

OpenStudy (anonymous):

Y=7